# Interval Selection in Sliding Windows

### v1

@@ 4.1.3 Sliding Window Algorithm
We call an even index $i$ good if there is an interval in $B_{2}^{1}$ that lies on $W_{i}W_{i+1}$. We consider the runs of $\mathcal{A}$ on pairs of windows $W_{i}W_{i+1}$ where $i$ is even. Then, we find a solution of size:
@@
Using the identity $X=\sum_{i=1}^{\ell}x_{i}$ and that $|B^{1}_{2}|$ is the number of good windows and proceeding as in the proof of Lemma 3, we obtain:
$S$ $\geq\frac{X+|B^{1}_{2}|}{2}\geq\frac{2-\varepsilon}{2(1+\beta)}\ell+\frac{1}{6}|\mathcal{CP}(B)|\geq\dots\geq\frac{7-3\varepsilon}{6(1+\beta)}\ell\ .$
@@
$\frac{7-3\varepsilon}{6(1+\beta)}\ell\ .$
@@
$\frac{\ell}{1+\beta}=\frac{7-3\varepsilon}{6(1+\beta)}\ell\ ,$
which implies $\varepsilon=\frac{1}{3}$. The approximation factor thus is for any $\delta>0$:
$4+2\cdot\beta-\varepsilon+\delta=11/3+2\cdot\beta+\delta\ .$
@@
The proof of the approximation factor of the algorithm is shown to be tight in Appendix A, meaning that the algorithm does not beat $\frac{11}{3}$ approximation.
@@ 4.2 Space Lower Bound
    $I_{1}(i)=\begin{cases}\left[\frac{i}{3n},1+\frac{i}{3n}\right],&\text{if }X[i]=1\ ,\\ [-10-i,10+i],&\text{if }X[i]=0\ .\end{cases}$
    $I_{2}(i)=\begin{cases}\left[2-\frac{i}{3n},3-\frac{i}{3n}\right],&\text{if }X[i]=1\ ,\\ [-11-i,11+i],&\text{if }X[i]=0\ .\end{cases}$
@@
The total number of intervals added by the three players is $3n+2+2(J_{1}-1)=L+2(J_{1}-1)$. So, after Charlie’s execution $\mathcal{A}$, the incumbent window indeed consists of $\mathcal{I}$.
@@
Also, we have that any interval $I_{1}(i)$ with $i>J_{1}$ would block all the intervals $I_{3}(j)$ for $1\leq j\leq n+2(J_{1}-1)$ and $I_{J_{2}}$. So, an interval from $I_{1}(i)$ with $i>J_{1}$ can be included in an optimal set of size at most $2$ (either $\{I_{1}(i),I^{\prime}_{J_{2}}\}$ or $\{I_{1}(i),I_{2}(j)\}$ for some $j>J_{1}$). Similarly, $I_{2}(i)$ with $i>J_{2}$ can be included in an optimal set of size at most $2$. Furthermore, we can construct from Bob and Charlie’s input a solution of size at most $2$ (similar to the $\frac{3}{2}-\varepsilon$ insertion only lower bound construction of [10]). The size of an optimal solution is thus in this case $2$.
@@ A.1 Description of the Instance
We call the created windows $[x,x+1)$ with $x=3k+2$, for some integer $k$, good. Notice that the number of windows created by $\mathcal{CP}(A)$ is $\ell$ while the number of good windows is exactly $\ell/3$ because $\ell$ is divisible by 3.
@@
Figure 6: Illustration of a good window, where the vertical black lines depict its boundaries. Black intervals represent streams $A_{2},A_{3}$. Blue intervals represent the stream $B$. The upper green intervals belong to $C_{1}$. The bottom green intervals belong to $C_{2}$. Stream $A_{1}$ is omitted to keep the illustration simple – these intervals are responsible for creating the windows of $\mathcal{CP}(A)$).
@@
      This stream is responsible for creating the windows $[x,x+1)$ for any integer $x\in[\ell]\setminus\{1,\ell\}$ and windows $(-\infty,2),[\ell,\infty)$.
@@
      This stream inserts intervals inside the windows $[x,x+1)$. Notice that the interval $[x+0.5,x+0.54]\in A_{2}$ is completely inside the window $[x,x+1)$ and replaces the interval $[x+0.1,x+1]\in A_{1}$ as both the leftmost and the rightmost intervals of the window.
@@
      This stream inserts intervals that intersect the boundaries of windows created by $A_{1}$. The stream contributes to the size of the optimal solution of the overall stream $ABC$.
@@
    If we execute stream $B$ immediately after stream $A$, the windows created in $A$ will remain unchanged. In $B$, we only add items inside or intersecting the good windows. Consider therefore a good window $W=[x,x+1)$.
    The intervals $[x-0.1,x+0.26],[x+0.9,x+1.1]$ are intervals crossing the boundary of $W$. They completely include the boundary intervals of $A_{3}$ (i.e $[x-0.05,x+0.05]$ from the previous window and $[x+0.95,x+1.05]$). The interval $[x+0.53,x+0.71]\in B$ is an interval inside $W$ that intersects the interval $[x+0.5,x+0.54]\in A_{2}$.
@@
    The stream $C$ is divided into $C_{1},C_{2}$ and only adds intervals completely included within good windows. Consider therefore a good window $W=[x,x+1)$.
@@
      The purpose of the stream $C_{1}$ is to create new windows from the windows of the run $\mathcal{CP}(AB)$. The new windows created inside $W$ are $[x,x+0.3),[x+0.3,x+0.75)$ and $[x+0.75,x+1)$.
@@
      The stream $C_{2}$ contributes to $|OPT\cap C|$ of our instance, where $OPT$ is an independent set of optimal size inside the stream $ABC$. The interval $[x+0.06,x+0.15]\in C_{2}$ intersects the interval $[x-0.1,x+0.26]\in B$, but it does not intersect the interval $[x-0.05,x+0.05]\in A_{3}$. The interval $[x+0.9,x+0.94]\in C_{2}$ has similar properties. The intervals $[x+0.25,x+0.35]\in C_{2}$ and $[x+0.7,x+0.8]\in C_{2}$ intersect with the boundaries of the windows of $\mathcal{CP}(C_{1})$, so they will not be saved by the algorithm. Lastly, the interval $[x+0.55,x+0.6]\in C_{2}$ does not intersect with $[x+0.5,x+0.54]\in A_{2}$, but it intersects with $[x+0.53,x+0.71]\in B$.
@@ A.2 Analysis of the Instance
Notice that after the run of stream $A$, we have created the windows $[x,x+1)$ for $x\in[\ell]$ and windows $(-\infty,2),[\ell,\infty)$.
Now, we consider a good window $W=[x,x+1)$. The saved interval of $A$ inside $W$ is $[x+0.5,x+0.54]$.
When the stream $B$ arrives, the intervals $[x-0.1,x+0.26]$ and $[x+0.9,x+1.1]$ cross the boundaries of $W$. The interval $[x+0.53,x+0.71]\in B$ intersects with the interval $[x+0.5,x+0.54]\in A$, so only the rightmost of the window is changed after the stream $B$ is processed.
Since no new windows are created by $B$, we can argue that $\mathcal{CP}(AB)=\ell$ (the number of windows created by $\mathcal{CP}(A)$). Furthermore, all the intervals of $B$ are pairwise independent so that $\mathcal{CP}(B)=|B|=\ell$, hence proving the required lemma.
@@
Recall that all intervals in $B$ and $C$ are inserted only into good windows or at the boundary of good windows. Let $[x,x+1)$ be a good window (i.e $x=3k+2$ for integer $k$). We will show that $|\mathcal{CP}(BC)\cap[x-1,x+2]|=3$ (the good window and its neighbouring windows).
After processing the $B$ stream, we have window boundaries at $x+0.26$, $x+0.71$ and $x+1.1$. Observe that all of the intervals of $C_{1}$ cross the window boundaries at $x+0.26$, and $x+0.71$, so they do not get saved by the run of $\mathcal{CP}$. Furthermore, we have that the interval $[x+0.25,x+0.35]\in C_{2}$ crosses the window boundary at $x+0.26$ while the interval $[x+0.7,x+0.8]\in C_{2}$ crosses the boundary at $x+0.71$, so these intervals also do not get saved.
When processing $C$, however, the intervals $[x+0.06,x+0.15],[x+0.55,x+0.6],[x+0.9,x+0.94]\in C_{2}$ get inserted into the solution. Additionally, they do not change the structure of the windows created by the run $\mathcal{CP}(B)$ (i.e they only modify the leftmost or the rightmost interval of each window created by $B$).
So, $|\mathcal{CP}(BC)\cap[x-1,x+2]|=3$. Because there are $\ell/3$ good windows and the intervals $[x-1,x+2]$ where $x=3k+2$ do not pairwise intersect, we have $|\mathcal{CP}(BC)|=\ell$ as required.
@@
First, observe that steps 1 and 2 of Algorithm 4 are run on substream $C$. Furthermore, since only good windows contain intervals in substream $C$, it suffices to explore how steps 1 and 2 act on good windows.
In each good window, the stream $C_{1}$ is responsible for creating the windows of the runs of Algorithm 4. Observe that the intervals $[x+0.25,x+0.35]\in C_{2}$ and $[x+0.7,x+0.8]\in C_{2}$ cross the boundaries of these windows and are thus not stored by the algorithm. In each good window, only the intervals $[x+0.06,x+0.15],[x+0.55,x+0.6],[x+0.9,x+0.94]$ of $C_{2}$ get memorized. Therefore, we obtain a solution of size $3$ for each good window. Overall, the obtained solution by the runs of steps 1 and 2 of Algorithm 4 is of size $3\cdot\frac{\ell}{3}=\ell$.

### v3

@@ 4.1.3 Sliding Window Algorithm
We call an even index $i$ good if there is an interval in $B_{2}^{1}$ that lies on $R_{i}R_{i+1}$. We consider the runs of $\mathcal{A}$ on pairs of regions $R_{i}R_{i+1}$ where $i$ is even. Then, we find a solution of size:
@@
Using the identity $X=\sum_{i=1}^{\ell}x_{i}$ and that $|B^{1}_{2}|$ is the number of good regions and proceeding as in the proof of Lemma 3, we obtain:
$S$ $\geq\frac{X+|B^{1}_{2}|}{2}\geq\frac{2-\varepsilon^{\prime}}{2(1+\beta)}\ell+\frac{1}{6}|\mathcal{CP}(B)|\geq\dots\geq\frac{7-3\varepsilon^{\prime}}{6(1+\beta)}\ell\ .$
@@
$\frac{7-3\varepsilon^{\prime}}{6(1+\beta)}\ell\ .$
@@
$\frac{\ell}{1+\beta}=\frac{7-3\varepsilon^{\prime}}{6(1+\beta)}\ell\ ,$
which implies $\varepsilon^{\prime}=\frac{1}{3}$. The approximation factor thus is for any $\delta>0$:
$4+2\cdot\beta-\varepsilon^{\prime}+\delta=11/3+2\cdot\beta+\delta\ .$
@@
The proof of the approximation factor of the algorithm is shown to be tight in Appendix A, meaning that the algorithm does not beat the approximation factor of $\frac{11}{3}$.
@@ 4.2 Space Lower Bound
    $I_{1}(i)=\begin{cases}\left[\frac{i}{3n},1+\frac{i}{3n}\right],&\text{if }X_{1}[i]=1\ ,\\ [-10-i,10+i],&\text{if }X_{1}[i]=0\ .\end{cases}$
    $I_{2}(i)=\begin{cases}\left[2-\frac{i}{3n},3-\frac{i}{3n}\right],&\text{if }X_{2}[i]=1\ ,\\ [-11-i,11+i],&\text{if }X_{2}[i]=0\ .\end{cases}$
@@
The total number of intervals added by the three players is $3n+2+2(J_{1}-1)=L+2(J_{1}-1)$. So, after Charlie’s execution $\mathcal{A}$, the incumbent region indeed consists of $\mathcal{I}$.
@@
Also, we have that any interval $I_{1}(i)$ with $i>J_{1}$ would block all the intervals $I_{3}(j)$ for $1\leq j\leq n+2(J_{1}-1)$ and $I_{J_{2}}$. So, an interval from $I_{1}(i)$ with $i>J_{1}$ can be included in an optimal set of size at most $2$ (either $\{I_{1}(i),I^{\prime}_{J_{2}}\}$ or $\{I_{1}(i),I_{2}(j)\}$ for some $j>J_{1}$). Similarly, $I_{2}(i)$ with $i>J_{2}$ can be included in an optimal set of size at most $2$. Furthermore, we can construct from Bob and Charlie’s input a solution of size at most $2$ (similar to the $\frac{3}{2}-\varepsilon$ lower bound construction of [10]). The size of an optimal solution is thus in this case $2$.
@@ 5 Conclusion
Acknowledgement
We are grateful to Sang-Sub Kim for identifying that the clean-up routine in the smooth histogram method used in an earlier version of this paper did not, as claimed, establish properties S1 and S2. This issue has been addressed and corrected in the current version.
@@ A.1 Description of the Instance
We call the created regions $[x,x+1)$ with $x=3k+2$, for some integer $k$, good. Notice that the number of regions created by $\mathcal{CP}(A)$ is $\ell$ while the number of good regions is exactly $\ell/3$ because $\ell$ is divisible by 3.
@@
Figure 6: Illustration of a good region, where the vertical black lines depict its boundaries. Black intervals represent streams $A_{2},A_{3}$. Blue intervals represent the stream $B$. The upper green intervals belong to $C_{1}$. The bottom green intervals belong to $C_{2}$. Stream $A_{1}$ is omitted to keep the illustration simple – these intervals are responsible for creating the regions of $\mathcal{CP}(A)$).
@@
      This stream is responsible for creating the regions $[x,x+1)$ for any integer $x\in[\ell]\setminus\{1,\ell\}$ and regions $(-\infty,2),[\ell,\infty)$.
@@
      This stream inserts intervals inside the regions $[x,x+1)$. Notice that the interval $[x+0.5,x+0.54]\in A_{2}$ is completely inside the region $[x,x+1)$ and replaces the interval $[x+0.1,x+1]\in A_{1}$ as both the leftmost and the rightmost intervals of the region.
@@
      This stream inserts intervals that intersect the boundaries of regions created by $A_{1}$. The stream contributes to the size of the optimal solution of the overall stream $ABC$.
@@
    If we execute stream $B$ immediately after stream $A$, the regions created in $A$ will remain unchanged. In $B$, we only add items inside or intersecting the good regions. Consider therefore a good region $R=[x,x+1)$.
    The intervals $[x-0.1,x+0.26],[x+0.9,x+1.1]$ are intervals crossing the boundary of $R$. They completely include the boundary intervals of $A_{3}$ (i.e $[x-0.05,x+0.05]$ from the previous region and $[x+0.95,x+1.05]$). The interval $[x+0.53,x+0.71]\in B$ is an interval inside $R$ that intersects the interval $[x+0.5,x+0.54]\in A_{2}$.
@@
    The stream $C$ is divided into $C_{1},C_{2}$ and only adds intervals completely included within good regions. Consider therefore a good region $R=[x,x+1)$.
@@
      The purpose of the stream $C_{1}$ is to create new regions from the regions of the run $\mathcal{CP}(AB)$. The new regions created inside $R$ are $[x,x+0.3),[x+0.3,x+0.75)$ and $[x+0.75,x+1)$.
@@
      The stream $C_{2}$ contributes to $|OPT\cap C|$ of our instance, where $OPT$ is an independent set of optimal size inside the stream $ABC$. The interval $[x+0.06,x+0.15]\in C_{2}$ intersects the interval $[x-0.1,x+0.26]\in B$, but it does not intersect the interval $[x-0.05,x+0.05]\in A_{3}$. The interval $[x+0.9,x+0.94]\in C_{2}$ has similar properties. The intervals $[x+0.25,x+0.35]\in C_{2}$ and $[x+0.7,x+0.8]\in C_{2}$ intersect with the boundaries of the regions of $\mathcal{CP}(C_{1})$, so they will not be saved by the algorithm. Lastly, the interval $[x+0.55,x+0.6]\in C_{2}$ does not intersect with $[x+0.5,x+0.54]\in A_{2}$, but it intersects with $[x+0.53,x+0.71]\in B$.
@@ A.2 Analysis of the Instance
Notice that after the run of stream $A$, we have created the regions $[x,x+1)$ for $x\in[\ell]$ and regions $(-\infty,2),[\ell,\infty)$.
Now, we consider a good region $R=[x,x+1)$. The saved interval of $A$ inside $R$ is $[x+0.5,x+0.54]$.
When the stream $B$ arrives, the intervals $[x-0.1,x+0.26]$ and $[x+0.9,x+1.1]$ cross the boundaries of $R$. The interval $[x+0.53,x+0.71]\in B$ intersects with the interval $[x+0.5,x+0.54]\in A$, so only the rightmost of the region is changed after the stream $B$ is processed.
Since no new regions are created by $B$, we can argue that $\mathcal{CP}(AB)=\ell$ (the number of regions created by $\mathcal{CP}(A)$). Furthermore, all the intervals of $B$ are pairwise independent so that $\mathcal{CP}(B)=|B|=\ell$, hence proving the required lemma.
@@
Recall that all intervals in $B$ and $C$ are inserted only into good regions or at the boundary of good regions. Let $[x,x+1)$ be a good region (i.e $x=3k+2$ for integer $k$). We will show that $|\mathcal{CP}(BC)\cap[x-1,x+2]|=3$ (the good region and its neighbouring regions).
After processing the $B$ stream, we have region boundaries at $x+0.26$, $x+0.71$ and $x+1.1$. Observe that all of the intervals of $C_{1}$ cross the region boundaries at $x+0.26$, and $x+0.71$, so they do not get saved by the run of $\mathcal{CP}$. Furthermore, we have that the interval $[x+0.25,x+0.35]\in C_{2}$ crosses the region boundary at $x+0.26$ while the interval $[x+0.7,x+0.8]\in C_{2}$ crosses the boundary at $x+0.71$, so these intervals also do not get saved.
When processing $C$, however, the intervals $[x+0.06,x+0.15],[x+0.55,x+0.6],[x+0.9,x+0.94]\in C_{2}$ get inserted into the solution. Additionally, they do not change the structure of the regions created by the run $\mathcal{CP}(B)$ (i.e they only modify the leftmost or the rightmost interval of each region created by $B$).
So, $|\mathcal{CP}(BC)\cap[x-1,x+2]|=3$. Because there are $\ell/3$ good regions and the intervals $[x-1,x+2]$ where $x=3k+2$ do not pairwise intersect, we have $|\mathcal{CP}(BC)|=\ell$ as required.
@@
First, observe that steps 1 and 2 of Algorithm 4 are run on substream $C$. Furthermore, since only good regions contain intervals in substream $C$, it suffices to explore how steps 1 and 2 act on good regions.
In each good region, the stream $C_{1}$ is responsible for creating the regions of the runs of Algorithm 4. Observe that the intervals $[x+0.25,x+0.35]\in C_{2}$ and $[x+0.7,x+0.8]\in C_{2}$ cross the boundaries of these regions and are thus not stored by the algorithm. In each good region, only the intervals $[x+0.06,x+0.15],[x+0.55,x+0.6],[x+0.9,x+0.94]$ of $C_{2}$ get memorized. Therefore, we obtain a solution of size $3$ for each good region. Overall, the obtained solution by the runs of steps 1 and 2 of Algorithm 4 is of size $3\cdot\frac{\ell}{3}=\ell$.